# Empowering Molecule Discovery for Molecule-Caption Translation with Large Language Models: A ChatGPT Perspective

### v1

@@ 4.2 Performance Comparison of Molecule-Caption Translation
We present the results of each sub-task within the molecule-caption translation task, incorporating both quantitative analysis and detailed examples for comparison. In addition, Figures 7, 5 and 4 illustrate specific examples that demonstrate the differences among various models, providing a visual understanding of their performance.
4.2.1 Molecule Captioning (Mol2Cap)
@@
Given a molecule’s SMILES representation, Mol2Cap aims to generate natural language for describing the molecule to enable humans to understand molecular structure, properties, and functionalities in a more efficient and cost-effective manner. Table 1 illustrates the performance comparison of 10-shot MolReGPT (GPT-3.5-turbo) with other advanced methods for the Mol2Cap task, offering an overview of the results. Notably, our method can achieve comparable ROUGE scores to MolT5-base and T5-base while surpassing all selected baselines in the remaining metrics without being fine-tuned on ChEBI-20 dataset. Furthermore, we obtain the following observations.
Table 1: The performance of molecule captioning on ChEBI-20 dataset. Experimental results for Transformer, T5-base, and MolT5-base are retrieved from  (Edwards et al., 2022). The best scores are in bold, and the second-best scores are underlined.
[…]
Transformer (Edwards et al., 2022) 0.061 0.027 0.204 0.087 0.186 0.114 0.057
T5-base (Edwards et al., 2022) 0.511 0.423 0.607 0.451 0.550 0.539 0.523
MolT5-base (Edwards et al., 2022) 0.540 0.457 0.634 0.485 0.578 0.569 0.547
GPT-3.5-turbo (zero-shot) 0.103 0.050 0.261 0.088 0.204 0.161 0.352
Table 2: Text-based molecule generation results on CheBI-20. Experimental results for Transformer, T5-base, and MolT5-base are retrieved from (Edwards et al., 2022). The best scores are in bold, and the second-best scores are underlined.
[…]
Transformer (Edwards et al., 2022) 0.499 0.000 57.66 0.480 0.320 0.217 11.32 0.277 0.906
T5-base (Edwards et al., 2022) 0.762 0.069 24.950 0.731 0.605 0.545 2.48 0.499 0.660
MolT5-base (Edwards et al., 2022) 0.769 0.081 24.458 0.721 0.588 0.529 2.18 0.496 0.772
@@
First, GPT-3.5-turbo is not explicitly trained or fine-tuned for molecule-caption translation tasks so it has poor zero-shot performance. However, with the instruction of 10-shot MolReGPT, GPT-3.5-turbo achieves significantly improved results that gain an improvement of 60% to the zero-shot case and 2.4% to MolT5-base under the Text2Mol metric, indicating that our proposed method can teach ChatGPT to effectively learn the Mol2Cap task from the system prompt.
@@
Second, limited by the number of examples, MolReGPT only gains limited insights from the distribution of molecule captions. The model’s predictions for captions heavily rely on its internal factual knowledge and the contextual information provided by the system prompt, which means that common patterns may not be as apparent and can not be captured from the selected $n$ examples. As a result, although our 10-shot MolReGPT achieves a 0.560 Text2Mol score, which is higher than MolT5’s 0.547, MolReGPT in turn gets lower ROUGE scores compared to MolT5. However, it is crucial to note that the captions generated by 10-shot MolReGPT with lower ROUGE scores are not entirely incorrect. In fact, the highest Text2Mol score serves as a reliable indicator of the generation quality and highlights the better relevance between the generated molecules and the molecule captions.
Figure 4 lists examples of molecule captioning results to compare the performance among different models. From the given examples, we note that our MolReGPT can generate captions that contain key information about the input molecule. And more importantly, the generated captions are better in grammar and easy for humans to understand.
@@
Figure 4: Examples of molecule captions generated by different models, where SMILES strings are converted to molecule graphs for better visualization. Based on the same input molecule graph, our MolReGPT can generate accurate and natural captions to describe the structure, properties, and even the functions of the molecule. In contrast, Transformer generates meaningless captions that are far from the ground truth. Captions generated by MolT5-base seem better but still have many typo errors.
4.2.2 Text-based Molecule Generation (Cap2Mol)
Given a natural language description (e.g., properties and functional groups), the goal of Cap2Mol is to generate the corresponding molecule (i.e., SMILES string) for molecule discovery. Results of the text-based molecule generation task are presented in Table 2. Comparing all these baselines, 10-shot MolReGPT significantly enhances the capabilities of GPT-3.5-turbo, leading to the best overall performance. In molecular evaluation metrics like MACCS FTS, RDK FTS, and Morgan FTS, MolReGPT helps GPT-3.5-turbo gain a significant 15% increase in Text2Mol score compared to MolT5-base. Considering the molecule fingerprints scores, our 10-shot MolReGPT also gets an average of 18% improvement compared to MolT5-base. Besides, MolReGPT also achieves the highest exact match score, generating 13.9% molecules that are completely correct to the ground truth. Remarkably, these impressive results are achieved without additional training or fine-tuning steps.
Furthermore, it is worth noting that the original weights of T5 are primarily for natural language generation, which means it has to be fine-tuned separately to fit the two sub-tasks in this study. Unfortunately, MolT5 does not tackle this issue, as it continues to treat the two sub-tasks of the molecule-caption translation task as separate tasks. Switching between the two sub-tasks in MolT5 requires using a different model class and reloading the weights, which makes it technically inefficient. Besides, treating these sub-tasks as independent overlooks the potential knowledge transfer between them. In contrast, MolReGPT enables a single foundation LLMs to solve both the two sub-tasks simultaneously, providing a comprehensive solution for LLMs to address molecule-related tasks.
Figure 5 lists examples of text-based molecule generation results to compare the performance among different models. From the given examples, it is clear that our MolReGPT can generate structures more similar to the ground truth.
@@ 4.2.2 Text-based Molecule Generation (Cap2Mol)
Figure 5: Examples of molecules generated by different models, where SMILES strings are converted to molecule graphs for better visualization. Based on the same input caption, our MolReGPT can generate accurate molecule graphs described by the caption. In contrast, Transformer generates quite different molecules compared to the ground truth. Compared to Transformer, molecules generated by MolT5-base are closer to the ground truth but still miss so many details.
4.3 Ablation Study
In addition to the experiments above, we also perform ablation studies to analyze the critical factors that influence the performance of MolReGPT. We first examine the impact of different retrieval strategies employed for retrieving n-shot examples. Subsequently, we investigate the influence of the number of selected examples, denoted as $n$, including zero-shot results to ensure that GPT-3.5-turbo is not already trained to handle the molecule-caption translation task. These ablation studies shed light on the key aspects contributing to the performance of MolReGPT.
4.3.1 Impact of Retrieval Strategies
Retrieval strategies play a key role in guiding LLMs to perform molecule-caption translation tasks for MolReGPT. More similar examples are retrieved, and more valuable information could be contained for in-context few-shot molecule learning. For each sub-task in the molecule-caption translation task, we choose three retrieval strategies for comparison. The detailed results are shown in Table 3 and Table 4.
Molecule Captioning (Mol2Cap). In the Mol2Cap task, we compare the performance of three retrieval strategies: Random, BM25, and Morgan FTS (adopted in MolReGPT),. The Random strategy involves retrieving $n$ random examples, while BM25 applies a character-level BM25 algorithm to the molecule SMILES representations.
As shown in Table 3, among the three retrieval strategies, Morgan FTS shows the best performance with the same value of $n$, outperforming BM25 by 37% in the Text2Mol metric. Besides, the ROUGE-L score achieved by Morgan FTS is almost doubled compared to the Random or BM25 retrieval strategies. The use of Morgan FTS with Dice similarity shows a better estimation of the structural similarity between molecules by comparing unique structural features like functional groups. These features are usually revealed in molecule captions with detailed descriptions. In this case, retrieving similar molecules by Morgan FTS could effectively guide the LLM to learn the associations between molecule structures and caption descriptions, resulting in more accurate and desired outputs.
Text-based Molecule Generation (Cap2Mol). In the Cap2Mol task, we also employ three retrieval strategies: Random, SentenceBert, and BM25 (adopted in MolReGPT). The Random strategy still retrieves $n$ random examples, while SentenceBert encodes captions as numerical vectors to compute their semantic similarity.
As shown in Table 4, we find that BM25 is better in the Cap2Mol task, despite the fact that SentenceBert has outperformed BM25 in many classical NLP text retrieval datasets. When $n$ changes from 1 to 10, n-shot BM25 always achieves better BLEU, Exact Match, Levenshtein, and fingerprints scores than n-shot SentenceBert. As shown in the input caption (stage#1) of Figure 2, the input molecule captions tend to use phrases with dashes (-) like ”2-methylphenyl” to connect the structure details of the molecule. Understanding such phrases plays a crucial role in generating correct molecule structures. In this case, retrieving similar texts while precisely matching these details significantly contributes to performance improvement. In contrast, SentenceBert, as a neural method, encodes an entire caption into a 1-D embedding vector, focusing more on semantic similarity rather than specific details. Consequently, BM25 is chosen as the retrieval strategy of MolReGPT in text-based molecule generation.
All in all, in both sub-tasks, compared to random selection, the other retrieval strategies used in this paper can help improve n-shot generation results. These strategies enhance the overall performance metrics, underscoring the importance of thoughtful retrieval strategy design for achieving performance improvement in MolReGPT.
4.3.2 Impact of the Number of Examples for In-Context Learning
Table 3: N-shot Molecule Captioning results on ChEBI-20 dataset. The best scores are in bold, and the second-best scores are underlined.
@@ 4.3.2 Impact of the Number of Examples for In-Context Learning
Table 4: N-shot Molecule Generation results on ChEBI-20 dataset. The best scores are in bold, and the second-best scores are underlined.
@@
Zero-shot Performance. In the zero-shot scenario, where no extra examples are included in the prompt for guiding LLMs for learning molecule-caption translation tasks, we utilize two special spans, ‘[MOLECULE_MASK]’ and ‘[CAPTION_MASK]’, to inform the LLMs of the desired output format, as shown in Figure 6 (in Appendix). In this case, the output of the LLMs can be conveniently filtered and further processed to satisfy the desired output specifications for molecule discovery.
@@
After analyzing the zero-shot results of GPT-3.5-turbo in Tables 3 and 4, we can observe that OpenAI did include SMILES strings in their training corpus because it can generate basically valid SMILES representations of molecules based on zero-shot prompts, achieving a 0.802 validity score and a 0.479 Text2Mol score in molecule generation. However, it is important to note that the zero-shot results exhibit a performance level similar to a vanilla Transformer model. This observation provides evidence that GPT-3.5-turbo is not specifically trained on the molecule-caption translation task, thereby alleviating concerns regarding potential information leakage.
Few-shot Performance. Table 3 and Table 4 list the comprehensive details of the experimental results, while Figure 8 (a) and Figure 8 (b) illustrate how the Text2Mol score changes when the number of examples increases.
@@
Normally, the performance should improve as the number of examples, denoted as $n$, increases, as more examples provide additional knowledge for the task at hand. However, due to the input length limitation of LLMs, it is impossible to contain a large number of examples in the system prompt. Therefore, for few-shot scenarios, we choose four different values 1, 2, 5, and 10.
@@
Tables 3 and 4 illustrate that performance generally improves as n increases in the system prompt through in-context learning. Significant performance enhancements are observed as n changes from 0 to 10. Taking Morgan FTS and BM25 as examples, in caption generation, we see remarkable increases from 0.050 to 0.482, 0.204 to 0.543, and 0.352 to 0.560 in BLEU-4, ROUGE-L, and Text2Mol scores, respectively. Besides, BM25 improves molecule generation from 0.489 to 0.790 in the BLEU score and 0.479 to 0.571 in the Text2Mol score.
@@
Besides, it is also interesting to notice that when $n$ increases from 5 to 10, the Text2Mol metrics almost keep the same. This could be the problem of the maximum input length limitation of LLMs. To fit the input length limitation, we would remove the longest examples to degrade the n-shot generation to (n-1)-shot generation. As $n$ increases, there is a higher possibility of exceeding the input length limitation. In this case, unless the maximum input length of the LLM is expanded, the performance will finally converge when $n$ continues to grow.
@@ 5 Conclusion
In this work, we propose MolReGPT, a general retrieval-based prompt paradigm that empowers molecule discovery with LLMs like ChatGPT under In-Context Few-Shot Molecule Learning. MolReGPT leverages the molecular similarity principle to retrieve examples from a local database, guiding LLMs in generating n-shot outputs without fine-tuning. Our method is focused and evaluated on the task of molecule-caption translation, including molecule captioning (Mol2Cap) and text-based molecule generation (Cap2Mol). Specifically, BM25 is applied to retrieve similar molecule captions, while Morgan Fingerprints and Dice similarity are adopted to retrieve similar molecules. Experimental results show that our proposed MolReGPT can empower ChatGPT to achieve 0.560 and 0.571 Text2Mol scores in molecule captioning and molecule generation, respectively. The performance surpasses fine-tuned models like MolT5-base in both molecule understanding and text-based molecule generation aspects and is even comparable to the fine-tuned MolT5-large. To conclude, MolReGPT provides a novel and versatile paradigm to deploy LLMs in molecule discovery through in-context learning, which greatly reduces the cost of domain transfer and explores the potential of LLMs in molecule discovery.
Appendix A Appendix
Figure 6: System Prompt for zero-shot Molecule-Caption translation. The main structure of zero-shot prompts is almost the same as that of few-shot prompts. The main difference lies in that the Example part in few-shot prompts is changed to Task Format to pre-define the input and output format. To avoid information leaks, we use ”[CAPTION_MASK]” and ”[MOLECULE_MASK]” to denote the position of captions and molecules.
Figure 7: Illustrations of molecule graphs generated by MolT5 and our MolReGPT, given customized inputs. Notably, the key points in Example 1 highlight the five benzene rings and hydrophobic groups in the structure, which are correctly generated by our MolReGPT. In contrast, the results of MolT5 generate the incorrect number of benzene rings and contain a few hydrophilic groups. In example 2, both generations give the correct number of benzene rings, while MolReGPT generates more hydrophilic groups, which are closer to our input caption.
(a) Text2Mol metric comparison of caption retrieval strategies with respect to the change of the number of selected examples in the Cap2Mol task.
(b) Text2Mol metric comparison of molecule retrieval strategies with respect to the change of the number of selected examples in the Mol2Cap task.
Figure 8: The trend of the Text2Mol metric with respect to the number of examples (i.e., $n$). Basically, as $n$ increases, the n-shot performance is also improved. However, when $n$ increases from 1 to 2, we see a clear performance drop in molecule generation, which is possibly the reason that the noise brought by the added examples exceeds the information gain they could bring. Besides, in caption generation, we see a remarkable increase by comparing Morgan Fingerprints to other retrieval strategies, showing the superiority of Morgan Fingerprints-based molecule retrieval. It is also interesting to notice that when $n$ increases from 5 to 10, the Text2Mol metrics almost keep the same. This is the problem of the maximum input length limitation of LLMs. To fit the input length limitation, we would remove the longest examples to degrade the n-shot generation to (n-1)-shot generation. As $n$ increases, there is a higher possibility of exceeding the input length limitation. In this case, unless the maximum input length of the LLMs is expanded, the performance will finally converge when $n$ continues to grow.

### v2

@@ 4.2 Performance Comparison
We present the results of the molecule-caption translation task, incorporating both quantitative analysis and detailed examples for comparison.
@@
Molecule Captioning (Mol2Cap). Table I illustrates the performance comparison of 10-shot MolReGPT (GPT-3.5-turbo) and 10-shot MolReGPT (GPT-4-0314) with the baseline models for Mol2Cap task, offering an overview of the results. Notably, our method can achieve better BLEU scores and comparable ROUGE scores to MolT5-base when using GPT-3.5-turbo as the backend model. Meanwhile, when using GPT-4-0314 as the backend model, MolReGPT can obtain better BLEU and Text2Mol scores than MolT5-large without being fine-tuned on ChEBI-20 dataset. Furthermore, we obtain the following observations:
@@
  • With the instruction of 10-shot MolReGPT, GPT-3.5-turbo achieves significantly improved results that gain an improvement of 59% to the zero-shot case and 2.4% to MolT5-base under the Text2Mol metric, which indicates that our proposed method can teach ChatGPT to effectively learn the Mol2Cap task via in-context learning.
@@
  • Restricted by the number of examples, MolReGPT only gains limited insights from the distribution of molecule captions (e.g., vocabulary and grammar). The LLMs’ predictions for captions heavily rely on their internal factual knowledge and the contextual information provided by the system prompt. Thus, such vocabulary and grammar patterns may not be as apparent and can not be captured from the selected $n$ examples. As a result, although our 10-shot MolReGPT (GPT-4-0314) achieves a 0.593 Text2Mol score, which is higher than MolT5-large’s 0.554. MolReGPT, in turn, gets lower ROUGE scores compared to MolT5. However, it is crucial to note that captions generated by 10-shot MolReGPT with lower ROUGE scores are not entirely incorrect. In fact, the highest Text2Mol score serves as a reliable indicator of the generation quality and highlights the better relevance between the generated molecules and the molecule captions.
Considering the model agnosticism, MolReGPT increases the caption generation performance consistently across the three different models. This is particularly evident with Llama-2-7B, as it possesses a different model structure. As depicted in Table I, Llama-2-7B achieves 0.409 BLEU-4 score and 0.466 Text2Mol score without further fine-tuning, which is close to the performance of T5-base. This denotes that MolReGPT is not dependent on specific models but is a general method that could be applicable to a broad spectrum of LLMs.
@@
Figure 4: Examples of molecule captions generated by different models, where SMILES strings are converted to molecule graphs for better visualization. Based on the same input molecule graph, our MolReGPT (10-shot GPT-4-0314) can generate accurate and natural captions to describe the structure, properties, and even the functions of the molecule. In contrast, Transformer generates meaningless captions that are far from the ground truth. Captions generated by MolT5-large seem better but still have some typo errors.
Figure 4 compares our predicted captions with the ground truth. It is clear that our prediction is quite close to the ground truth, stating accurate details of molecule structures and giving solid predictions of molecule properties. However, it can also be seen that our results have some slight differences in the narrative order, which might influence the translation scores like BLEU and ROUGE. In Contrast, although MolT5-large achieves higher ROUGE scores, the generated captions still make many fact and typo errors, contributing to a lower Text2Mol score. Besides, Transformer could hardly generate correct or valid results. For molecule 2 and 3, it even generates the same results, which means that the model is not sensitive to the input.
Text-based Molecule Generation (Cap2Mol).
@@
Figure 5: Examples of molecules generated by different models, where SMILES strings are converted to molecule graphs for better visualization. Based on the same input caption, our MolReGPT (10-shot GPT-4-0314) can generate accurate molecule graphs described by the caption. In contrast, Transformer generates quite different molecules compared to the ground truth. Compared to Transformer, molecules generated by MolT5-large are closer to the ground truth but still miss so many details.
Results of the text-based molecule generation task are presented in Table II. Comparing all these baselines, 10-shot MolReGPT significantly enhances the capabilities of GPT-3.5-turbo and GPT-4-0314, leading to improved overall performance. In the Text2Mol metric, MolReGPT helps GPT-3.5-turbo and GPT-4-0314 gain a significant 15% and 20% increase, respectively, compared to MolT5-base. What’s more, GPT-4-0314 even achieves a 7% improvement compared to MolT5-large. Considering the fingerprint scores, our 10-shot MolReGPT (GPT-4-0314) even gets an average of 8.1% improvement compared to MolT5-large. More importantly, MolReGPT also significantly enhances the molecule generation capabilities of Llama-2-7B, elevating the BELU score from 0.104 to 0.693. Given that Llama-2-7B was trained on a distinct corpus and possesses a different model architecture, it could further demonstrate the model-agnostic nature of our approach.
Figure 5 compares our predicted molecules with the ground truth. It can be seen that our generated molecules are quite close to the ground truth in the molecular configuration. For molecule 1, both MolT5-large and MolReGPT generate the exact correct molecule, but the 2D graph is slightly different to the ground truth due to the sequence order of the SMILES string. For molecule 2, MolT5-large misses several key structures, while MolReGPT generates the correct representation. For molecule 3, MolT5-large fails to generate the correct configuration, while MolReGPT generates the correct chemical bonds but misses the correct number of carbon atoms.
Figure 6: Illustrations of molecule graphs generated by MolT5 and our MolReGPT (GPT-3.5-turbo), given customized inputs. Notably, the key points in Example 1 highlight the five benzene rings and hydrophobic groups in the structure, which are correctly generated by our MolReGPT. In contrast, the results of MolT5 generate the incorrect number of benzene rings and contain a few hydrophilic groups. In example 2, both generations give the correct number of benzene rings, while MolReGPT generates more hydrophilic groups, which are closer to our input caption.
Besides, in Figure 6, we propose a practical scenario where a scientist aims to obtain molecules with desired structures and properties. In the past, scientist needs to apply domain knowledge to figure out a possible molecule candidate. Then, experiments are required to verify its properties. Now, with the help of MolReGPT, the scientist could formalize the requirements via molecule captions and ask MolReGPT to generate the desired molecules. To be more specific, we list two examples here. Molecule 1 has five benzene rings and several hydrophobic groups as its unique pattern, while molecule 2 has two benzene rings and several hydrophilic groups. By counting the number of benzene rings, we can easily find that MolT5 generates 3 benzene rings in molecule 1, which is incorrect. Besides, the functional groups generated by MolT5 in both molecule 1 and molecule 2 also miss-match the requirements given in the captions. Remarkably, these impressive results are achieved without additional fine-tuning steps.
@@
TABLE III: N-shot Molecule Captioning results on ChEBI-20 dataset with the backend model, GPT-3.5-turbo. The best scores are in bold, and the second-best scores are underlined.
@@
TABLE IV: N-shot Molecule Generation results on ChEBI-20 dataset with the backend model, GPT-3.5-turbo. The best scores are in bold, and the second-best scores are underlined.
Furthermore, the original weights of T5 are primarily for natural language, which means it has to be fine-tuned separately to fit the two sub-tasks in this study. Unfortunately, MolT5 does not tackle this issue, as it continues to treat the two sub-tasks of the molecule-caption translation task separately. Switching between the two sub-tasks in MolT5 requires using a different model class and reloading the weights, making it technically inefficient. In contrast, MolReGPT enables a single foundation LLM to solve both the two sub-tasks simultaneously, providing a comprehensive solution for LLMs to address molecule-related tasks.
4.3 Ablation Study
In addition to the experiments above, we also perform ablation studies to analyze the critical factors that influence the performance of MolReGPT. Note that we select GPT-3.5-turbo as the backend model.
4.3.1 Impact of Retrieval Strategies.
Retrieval strategies play a key role in guiding LLMs to perform molecule-caption translation tasks for MolReGPT. More similar examples are retrieved, and more valuable information could be contained for In-Context Few-Shot Molecule Learning. For Mol2Cap and Cap2Mol, we choose three different retrieval strategies for comparison. The detailed results are shown in Table III and Table IV. We show that in both sub-tasks, compared to random selection, the other retrieval strategies used in this paper can help improve n-shot generation results. Thus, thoughtful selection of retrieval strategies plays a key role in MolReGPT.
In Mol2Cap task, we compare the performance of three retrieval strategies: Random, BM25, and Morgan FTS (adopted in MolReGPT). The Random strategy involves retrieving $n$ random examples, while BM25 applies a character-level BM25 algorithm to the molecule SMILES representations. As shown in Table III, among the three strategies, Morgan FTS shows the best performance when $n$ is fixed, outperforming BM25 by 37% in the Text2Mol metric. Besides, the ROUGE-L score achieved by Morgan FTS is almost doubled compared to the Random or BM25 strategies. The use of Morgan FTS with Dice similarity shows a better estimation of the structural similarity between molecules by comparing unique structural features like functional groups. These features are usually revealed in molecule captions with detailed descriptions. In this case, retrieving similar molecules by Morgan FTS could effectively guide the LLM to learn the associations between molecule structures and caption descriptions, resulting in more accurate and desired outputs.
In Cap2Mol task, we also employ these retrieval strategies: Random, SentenceBert, and BM25 (adopted in MolReGPT). The Random strategy still retrieves $n$ random examples, while SentenceBert encodes captions as numerical vectors to compute their semantic similarity. As shown in Table IV, BM25 is the best one in the Cap2Mol task, despite the fact that SentenceBert has outperformed BM25 in many classical NLP text retrieval datasets. When $n$ changes from 1 to 10, BM25 always achieves better BLEU, Exact Match, Levenshtein, and fingerprint scores than SentenceBert. The input molecule captions tend to use phrases with dashes (-) like ”2-methylphenyl” to connect the structure details of the molecule. Understanding such phrases plays a crucial role in generating correct molecule structures. In this case, retrieving similar texts while precisely matching these details significantly contributes to performance improvement. In contrast, SentenceBert, as a neural method, encodes an entire caption into a 1-D embedding vector, focusing more on semantic similarity rather than specific details. Consequently, BM25 is better than SentenceBert in the Cap2Mol task.
4.3.2 Impact of Example Number for In-Context Learning.
@@ 4.3 Ablation Study
In the zero-shot scenario, where no extra examples are included in the prompt for guiding LLMs for learning the molecule-caption translation task, we utilize two special spans, ‘[MOLECULE_MASK]’ and ‘[CAPTION_MASK]’, to inform the LLMs of the desired output format.
@@
After analyzing the zero-shot results of GPT-3.5-turbo in Tables III and IV, we can observe that SMILES strings are included in its pre-training corpus because it can generate basically valid SMILES representations of molecules based on zero-shot prompts, achieving a 0.802 validity score and a 0.479 Text2Mol score in molecule generation. However, it is important to note that the zero-shot results exhibit a performance level similar to a vanilla Transformer model. This observation provides evidence that GPT-3.5-turbo is not specifically trained on ChEBI-20, thereby alleviating concerns regarding potential information leakage.
@@
For Few-shot Performance, Table III and Table IV list the comprehensive details of the experimental results.Normally, the performance improves as the number of examples, denoted as $n$, increases, as more examples provide additional knowledge for the task. However, due to the input length limitation of LLMs, it is impossible to contain a large number of examples in the system prompt. Therefore, for few-shot scenarios, we choose four different values 1, 2, 5, and 10.
@@
Tables III and IV illustrate that performance generally improves as $n$ increases through in-context learning. Significant performance enhancements are observed as n changes from 0 to 10. Taking Morgan FTS and BM25 as examples, in caption generation, we see remarkable increases from 0.050 to 0.482, 0.204 to 0.543, and 0.352 to 0.560 in BLEU-4, ROUGE-L, and Text2Mol scores, respectively. Besides, BM25 improves molecule generation from 0.489 to 0.790 in the BLEU score and 0.479 to 0.571 in the Text2Mol score.
@@
Besides, it is worth noticing that when $n$ increases from 5 to 10, the Text2Mol metrics almost keep the same, which can be the problem of the maximum input length limitation of LLMs. To fit the input length limitation, we would remove the longest examples to degrade the n-shot to (n-1)-shot generation for re-queries. As $n$ increases, there is a higher possibility of exceeding the input length limitation. In this case, unless the maximum input length of the LLM is expanded, the performance will finally converge when $n$ continues to grow.
@@ 5 Conclusion
In this work, we propose MolReGPT, a general retrieval-based in-context learning paradigm that empowers molecule discovery with LLMs like ChatGPT via In-Context Few-Shot Molecule Learning. Our method is focused and evaluated on the task of molecule-caption translation, including molecule captioning (Mol2Cap) and text-based molecule generation (Cap2Mol). MolReGPT leverages the molecular similarity principle to retrieve examples from a local database, guiding LLMs to generate predictions without being fine-tuned. Specifically, BM25 caption retrieval is applied to obtain similar molecule captions, while Morgan Fingerprints and Dice similarity are adopted to retrieve similar molecules. Experimental results show that our proposed MolReGPT can empower ChatGPT to achieve 0.585 and 0.593 Text2Mol scores in Mol2Cap and Cap2Mol, respectively. Compared to MolT5-large, our MolReGPT equips GPT-4-0314 with the ability to achieve comparable performance in Mol2Cap task and even outperform MolT5-large in Cap2Mol task without any fine-tuning steps. To conclude, MolReGPT provides a novel and versatile paradigm to deploy LLMs in molecule discovery through in-context learning, which greatly reduces the cost of domain transfer and explores the potential of LLMs in molecule discovery.
6 Broader Implication & Future Directions
Our work has demonstrated the effectiveness of the MolReGPT framework across various LLMs. Compared to the existing methods, MolReGPT does not require additional pre-training or fine-tuning, yet it enables powerful LLMs like GPT-4 to achieve comparable and even superior performance. The integration of LLMs and biomolecular science represents a paradigm shift in molecule discovery. As illustrated in Figure 6, fine-tuned methods tend to fail to meet the requirements in customized inputs, while MolReGPT shows a better generalization potential. These unique features make our method more convenient and practical for chemists to adopt and use in their work. We hope that MolReGPT could enable a wider range of scientific researchers to leverage the power of LLMs in their work, ultimately contributing to the advancement of scientific research and discovery.
For future work, there are still several areas that require further exploration and improvement. Firstly, the performance of LLMs is intricately related to the prediction quality of MolReGPT. Thus, with the advancement of more powerful LLMs, our methods could potentially yield even better results in the future. Secondly, we could develop better retrieval algorithms that could help refine the context examples. For example, we could combine BM25 caption retrieval with chemical LLMs and apply Graph Neural Networks for molecular similarity to improve the retrieval quality. Lastly, we anticipate that our work will not only make LLMs more accessible for scientific researchers in the field of chemistry, thereby benefiting drug discovery, but also inspire the AI community to consider the alignment between molecular and text space.